If the consumption function is C = 100 + 0.8Y, planned investment spending is 400, government expenditure is also 400 and the net exports 100, calculate what is the equilibrium level of output? If planned investment falls by 100 and net exports by 50 what is going to be the impact on the final output?
1
Expert's answer
2020-12-10T13:44:07-0500

Equilibrium level of output Y=C+I+G+X, so Y=100+0.8Y+400+400+100; then 1000+0.8Y=Y, so Y=1000/(1-0.8)=5000.

After investment and net exports fall the final Y equals to

Y=100+0.8Y+300+400+50; Y=850/0.2=4250